40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 30 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  1146   Tue Apr 9 15:39:16 2013 taraNotesPMCAbout PMCs

Considerations for PMC design:

  1. Stiffness(Acoustic susceptibility) & heavy material: With heavier material, the pmc motion on the support becomes smaller.(RXA: please quantify with a formula)
  2. Filtering factor (Finesse/FSR/Cavity pole), g-factor: Filter out intensity noise around 10 MHz (RXA: please quantify with a formula)
  3. Design for thermal expansion cancellation between the spacer and the end cap: So that the PMC is less sensitive to ambient temperature
  4.  3 or 4 mirrors?  3 is polarization selective. For general lab use with power less than 1 W,  3 mirror design should be good. (RXA: I don't follow this logic at all)

RXA: In general, all of these considerations need some sort of quantitative detail. Make a DeBra Matrix so that we can evaluate. 

  1145   Mon Apr 8 17:44:20 2013 EvanNotesPMCAbout PMCs

[Rana, Tara, Evan, Eric, Nic] 

We are designing a PMC, to do that we should be able to answer some fundamental questions about a PMC.

Why do we want a PMC?

  • Intensity stabilization(is it?)
  • Filtering of beam profile
  • Alignment reference
  • Reference for waist size and position
  • Jitter suppression
  • Polarization filtering

What should we consider in the design of the cavity?

  • High finesse eventually causes the transmission to drop
  • Cavity g should be chosen so that higher-order PMC modes are well outside the bandwidth of the TEM00 mode
  • Finesse should be chosen so that laser intensity noise is suppressed below the shot noise limit at frequencies of interest
  • Number of mirrors - determines whether or not we get polarization filtering
    • In a 3-mirror cavity, when one polarization is resonant, the other will be antiresonant. In a 4-mirror cavity, both will resonate simultaneously (with small differences in transmission intensity and phase).

What should we consider in the design of the spacer?

  • Acoustic pickup (probably dominates over seismic pickup at frequencies we are interested in)
  • Steel or silica:
    • Longitudinal mode frequencies are comparable
    • Silica is harder to machine
  • Vacuum can?
  • Dimensions:
    • A shorter PMC will have higher vibrational modes
    • A longer PMC has more filtering
    • A spacer that is too thin will have easily excited bending modes

Other:

  • 15 MHz is too low for PDH modulation, because the bandwidth of the resulting loop is too low
  • For a resonant PD, we want a low Q (a few)
  • 3 mirror or 4 mirror design?

 

  1144   Mon Apr 8 15:15:06 2013 ranaHowToTempCtrlestimated beat frequency

Isn't heating up one cavity enough? The goal is to keep the beat frequency constant, so we need to control the differential length between the two cavities. The first cavity can sit still, the second one can be heated up (for enough cooling rate). We plan to use a frequency counter to change beat frequency to the servo's error signal for feedback to the second cavity. 

Anyway, I still have to check if both heaters are working or not.

Quote:

I would have guessed that you heat both cavities. Unless they are both at an elevated temperature, how can you control their individual temperatures?

The cooling rate (and thus the bandwidth for control) is determined by the steady state temperature. I would guess that each cavity needs to be at least 35 C in order to have some headroom.

 

  1143   Sat Apr 6 09:32:35 2013 ranaHowToTempCtrlestimated beat frequency

I would have guessed that you heat both cavities. Unless they are both at an elevated temperature, how can you control their individual temperatures?

The cooling rate (and thus the bandwidth for control) is determined by the steady state temperature. I would guess that each cavity needs to be at least 35 C in order to have some headroom.

  1142   Fri Apr 5 13:56:01 2013 taraNotesRefCavestimated beat frequency

Note:A test to check which cavity needs to be heated up.

1) when heaters (on the shields) are off, C3:PSL-VAC_CHMBRTMP =31.2

  • CAV1 is locked @ (slow out Coarse/fine) = 774/500
  • CAV2                 @ 769/500

2 when heater on CAV2(4V), C3:PSL-VAC_CHMBRTMP = 31.2

  • CAV1             @ 775/500
  • CAV2              @ 756/500

So to bring both cavity to be resonant at the same time, the heater on cav1 should be on.

I'm not sure if the thermometers on the shields are working or not, I'll check them.

Quote:

At current temperature, the estimated beat frequency will be ~ 60-100MHz. This is not so bad, since we can use 1811 to measure the beat signal and use PLL to extract the beat noise.

 

We will need to use thermal expansion to tune the beat frequency. So, as a start, I try to figure out the beat frequency, and how much we have to heat up the cavity. The heaters on each cavity is off, only the heater around the vac chamber is on (but the servo is off). 

Right now we have one laser locked to one cavity, but the beam path to the first cavity has a beam splitter that we can borrow the beam and direct it to the 2nd cavity. I realigned the beam to have both beams into both cavities. By adjusting the temperature control on the NPRO (slow signal), I can bring the beam to resonant in each cavity.

1st cav is resonant @ (334/398) and (150/398). The numbers correspond to coarse and fine knobs of the slow feedback to the laser.

2nd cav is resonant @ (154/398). (I'll come up with a better name to call the cavities)

The FSR is 4.07 GHz (for 1.45" long cavity). This means 334-154 = 184 clicks on the coarse knob equals to 4.07 GHz, or 22MHz per coarse click. Both cavities resonant at ~3-5 clicks apart. So the beat frequency is ~ 60-100MHz. This is quite good, at least we are not close to half FSR apart. The power required to tune the cavity length should not be that high.

 

The next thing to do is try to see which cavity we need to heat up in order to bring both cavities resonant frequency closer together.

 

  1141   Thu Apr 4 23:55:12 2013 EvanNotesPMCPMC Longitudinal Testing and Modal Analysis

 

These are plots of the sagging of the front and back mirrors as a function of the longitudinal positions of the mounting holes (these positions are measured from the back of the PMC). The first plot is a coarse search, and the second is more targeted toward a region of lower sagging.

I generated these plots by taking Tara's Comsol model of the PMC body, assigning fixed displacement to the three mounting holes, and assigning a body load to the PMC body equal to the weight of the steel. Then, I extracted the displacements of four points on the front edge and four points on the back edge of the PMC borehole (these edges are where the faces of the mirrors will make contact with the body). I then took some cross-products with these points in order to get the unit normals that would result when the mirrors are placed against the deformed body. I then compute the angle between the deformed unit normals and the undeformed unit normals to get the sag of the mirrors in radians.

I'm a bit uneasy about how precision is handled in the Comsol/Matlab combination used to generate these plots. The Comsol GUI has no problem reporting displacements all the way down to 10^-24 meters, but anything smaller than 10^-15 meters or so gets truncated to exactly 0 when the results are reported in Matlab. When propagated through to the sagging computation, this means any sagging smaller than 10^-8 radians or so also gets rounded to exactly 0. You can see in the second set of plots that there are large swaths of exactly the same light blue and periwinkle, which seems to indicate a low level of precision in the computation. There's probably some obvious Comsol/Matlab setting that I'm missing, but I haven't been able to find it so far.

Regardless, it appears there is an optimum range of hole placements for the PMC body: 10 cm for the front holes and 3 cm for the back holes, give or take a centimeter or so.

Quote:

I calculated some requirement for the beam jitter at the output of the PMC. A rough estimate shows that we need the angular stability at the PMC about half nano radian so that the frequency noise of the beam locked to the refcav is less than 10-2 Hz/rtHz.

==Background==

PMC also reduces beam jitters from the laser, so that the beam alignment to the cavity is kept centered. Since the laser is locked to the reference cavity, any misalignment of the input beam will cause the beam to sense the change of the cavity length.

So vibration that shakes the PMC will change the alignment of the output beam. With stiff material, the seismic induced deformation of the PMC will be reduced.

==Calculation==

  • calculate the ray tracing matrices from the PMC to the cavity. I assume that only the angle of the output beam changes due to PMC sagging, because of a long distance from the PMC to the refcav, with several mirrors in between. This gives me the position and the angle of the beam going to the cavity.
  • find out what is the change of the cavity length (dL), when the input beam is translated by dx, with angle theta.
  • convert displacement noise to frequency nosie (dL -> df), as a rough estimate I choose the requirement for df to be less than 10-2 Hz/rtHz (about the level of the estimated coating noise). This step is not really necessary, but I feel that it is easier to compare the noise in Hz/rtHz unit rather than m/rtHz.
  • The required angular stability at the PMC is ~ 0.5 nano rad. This number seems to be too strict. I will double check it.

==next==

Eavn is working on COMSOL to find out the angular tilt of the output beam due to PMC sagging. Optimum support points will be determined to minimize beam jitter due to seismic.

Attachment 1: mirrorsag_coarse.pdf
mirrorsag_coarse.pdf
Attachment 2: mirrorsag_fine.pdf
mirrorsag_fine.pdf
  1140   Thu Apr 4 20:46:03 2013 EvanDailyProgressElectronics EquipmentModified second TTFSS

I made some modifications to TTFSS box S0900371, so that it is more like the TTFSS box we are currently using for the south cavity. The modifications are as follows:

  • R1, R2, and R7 are now 47 Ω
  • R3 is now 100 Ω
  • R4 is now 390 Ω
  • R29 is now 5.1 kΩ, with an optional 6.8 nF series capacitor controlled by a switch on the side of the box. This duplicates the optional boost available on the other TTFSS.

I checked my soldering work with an LCR meter and everything seems fine, but I have not yet powered it up.

  1139   Thu Apr 4 18:45:38 2013 taraNotesDAQTemperature control updated

[peter, tara]The temperature servo for the chamber is back on, the current setup is at 31.2 C.

There was a problem with C3:PSL-VAC_CHAMBERTEMP channel, and I could not run the script for temperature control of the chamber. Peter helped me figure out what happened. It turned out that one of the parenthesis in the database file (cavities.db) was missing due to an accidental delete, and the name of the channel was too long (it was working before, I don't know why).

Anyway, the channel was renamed to C3:PSL-VAC_CHMBRTMP, in 1)cavities.db, 2)rcav_PID_2012_06_15.pl, and 3) medm screen for controlling the servo. The temperature servo is working again.

Quote:

We switched the temperature readout channels used for temp feedback control to improve the signal. The new signal is significantly smoother.

      The signals from 4 thermostats around the vacuum chamber were acquired through 4 channels, C3:PSL-RCAV_SENSE(1-4). These channels were then connected to DAQ. This made the signal noisy because the resolution of analog to digital converter was low. In order to fix that we use an analog circuit to sum and average the signals from 4 sensors then amplify it before sending to DAQ,C3:PSL-RCAV_TEMP, then calibrated it to C3:PSL-VAC_CHAMBERTEMP by comparing RCAV_TEMP [V] to RCAV_TEMPAVG[C] which is calibrated to deg C already. 

      CHAMBERTEMP = (RCAV_TEMPx-0.495) + 34.957

     

     We corrected the perl script (in SUN machine) used for thermal feedback on the heater jacket. Now the script is named rcav_PID_2012_06_15.pl, see wiki. The servo is now back on.

 

  1138   Thu Apr 4 11:44:14 2013 EvanNotesPMC270 Hz clamped PMC twisting mode

I ran another Comsol simulation with a simplified version of the PMC spacer. This time I put fixed constraints on two circular regions on the sides of the PMC near where it was clamped for the ringdown measurement. Comsol says the spacer has a mode where it twists about these clamp points, and the frequency of the mode is 270 Hz.

Attachment 1: pmc_clamp_twist_mode.jpg
pmc_clamp_twist_mode.jpg
  1137   Mon Apr 1 20:59:19 2013 taraNotesPMCPMC Longitudinal Testing and Modal Analysis

I calculated some requirement for the beam jitter at the output of the PMC. A rough estimate shows that we need the angular stability at the PMC about half nano radian so that the frequency noise of the beam locked to the refcav is less than 10-2 Hz/rtHz.

==Background==

PMC also reduces beam jitters from the laser, so that the beam alignment to the cavity is kept centered. Since the laser is locked to the reference cavity, any misalignment of the input beam will cause the beam to sense the change of the cavity length.

So vibration that shakes the PMC will change the alignment of the output beam. With stiff material, the seismic induced deformation of the PMC will be reduced.

==Calculation==

  • calculate the ray tracing matrices from the PMC to the cavity. I assume that only the angle of the output beam changes due to PMC sagging, because of a long distance from the PMC to the refcav, with several mirrors in between. This gives me the position and the angle of the beam going to the cavity.
  • find out what is the change of the cavity length (dL), when the input beam is translated by dx, with angle theta.
  • convert displacement noise to frequency nosie (dL -> df), as a rough estimate I choose the requirement for df to be less than 10-2 Hz/rtHz (about the level of the estimated coating noise). This step is not really necessary, but I feel that it is easier to compare the noise in Hz/rtHz unit rather than m/rtHz.
  • The required angular stability at the PMC is ~ 0.5 nano rad. This number seems to be too strict. I will double check it.

==next==

Eavn is working on COMSOL to find out the angular tilt of the output beam due to PMC sagging. Optimum support points will be determined to minimize beam jitter due to seismic.

  1136   Sun Mar 31 20:06:16 2013 taraNotesPMCPMC Longitudinal Testing and Modal Analysis

Quote:

I think the analytical formula in terms of rho is going to be (1.57/2*pi) * sqrt(E / rho * L^2), since the Roark formula is (1.57/2*pi) * sqrt(A* E * g / w * L^2) and the weight per unit length is w = m * g / L = rho * A * g. With your values for L, A, E, and rho, this gives f1 = 16 kHz. Since A does not appear in the analytical formula, this also explains why changing the area in the Comsol model doesn't change the frequency.

 

 

 good catch! Thanks. Then both analytical and FEA results are the same. So our COMSOL results for PMC should be valid, the first body for a stainless steel PMC, see psl:1131,at 16 kHz is reasonable.

  1135   Sun Mar 31 14:42:05 2013 EvanNotesPMCPMC Longitudinal Testing and Modal Analysis

I think the analytical formula in terms of rho is going to be (1.57/2*pi) * sqrt(E / rho * L^2), since the Roark formula is (1.57/2*pi) * sqrt(A* E * g / w * L^2) and the weight per unit length is w = m * g / L = rho * A * g. With your values for L, A, E, and rho, this gives f1 = 16 kHz. Since A does not appear in the analytical formula, this also explains why changing the area in the Comsol model doesn't change the frequency.

Quote:

I compared results between COMSOL and analytical solution. The first longitudinal mode from both results differ by an order of magnitude!!

Peter sent me a note from Dennis about PMC longitudinal mode calculation. Dennis mentioned about a book by Young&Roark (here), so I looked it up and see how to estimate body mode frequencies of a simple block/beam.  I tried a simple geometry, a 0.1x0.1x0.175 (m) block. According to the book, cf situation 7b, table16.1 page 771, the first longitudinal mode is

f1 = (1.57/2*pi) * sqrt ( AE/ rho*L^2), where A is the cross section area (0.1x0.1), rho is the mass density of the material (2202 kg/m^3, for SiO2), E is the Young's modulus (72 GPa), L is the length of the block ( I use L = 0.175/2 because 7b situation is a uniform bar vibrates along its longitudinal axis, with upper end fixed, lower end free. This is similar to a whole beam resonate freely on both end because its center will be fix. Thus, to use the formula for our case, we have to use half length of the beam).

The analytical solution gives f1 = 1.6 kHz ,while COMSOL result is ~ 16 kHz.

psl_log2.png

 It is very strange that, according to COMSOL simulation, when the cross sectional area of the block is changed to 0.01x0.01 m^2 instead of 0.1x0.1 m^2, the frequency of the longitudinal mode does not change that much (still close to 16kHz. However, from the analytical solution, the frequency should drop by a factor of 10 ( around 165 Hz).

I'm going to think about this a bit more, but at this point, I think my COMSOL model is not correct. Might be some kind of bdy conditions that I'm missing.

psl_log.png

 

 

 

 

  1134   Sun Mar 31 03:03:28 2013 taraNotesPMCPMC Longitudinal Testing and Modal Analysis

[see PSL:1135]

I compared results between COMSOL and analytical solution. The first longitudinal mode from both results are comparable.

Peter sent me a note from Dennis about PMC longitudinal mode calculation. Dennis mentioned about a book by Young&Roark (here), so I looked it up and see how to estimate body mode frequencies of a simple block/beam.  I tried a simple geometry, a 0.1x0.1x0.175 (m) block. According to the book, cf situation 7b, table16.1 page 771, the first longitudinal mode is

f1 = (1.57/2*pi) * sqrt ( E/ rho*L^2), ), rho is the mass density of the material (2202 kg/m^3, for SiO2), E is the Young's modulus (72 GPa), L is the length of the block ( I use L = 0.175/2 because 7b situation is a uniform bar vibrates along its longitudinal axis, with upper end fixed, lower end free. This is similar to a whole beam resonate freely on both end because its center will be fix. Thus, to use the formula for our case, we have to use half length of the beam).

The analytical solution and COMSOL give f1 ~ 16 kHz.

psl_log2.png

 It is very strange that, according to COMSOL simulation, when the cross sectional area of the block is changed to 0.01x0.01 m^2 instead of 0.1x0.1 m^2, the frequency of the longitudinal mode does not change that much (still close to 16kHz. However, from the analytical solution, the frequency should drop by a factor of 10 ( around 165 Hz).

I'm going to think about this a bit more, but at this point, I think my COMSOL model is not correct. Might be some kind of bdy conditions that I'm missing.

psl_log.png

 

 

 

  1133   Sat Mar 30 01:42:16 2013 taraDailyProgressElectronics EquipmentFrequency readout noise of PDH loop

It is indeed really high. We expect something in the order of ten nV/rtHz level (like what we had before, see PSL:781.). We are investigating it.

One of the possible causes is the back reflected beam. When I reduced the power input from 6mW to 1mW (with common/fast gain = 999/720, boost on), the error noise was down to ~ 10nV/rtHz up to 5 kHz, with bumps and peaks around 50kHz up to 100kHz. Our mode matching this time is not very good.

 I replaced the PBS for PDH locking to a bigger one, since the beam spot at that position was quite large. This got rid off some high frequency peaks and bumps. I still have to align the beam to minimize back reflection.

 

Quote:

Is this really the input noise of the PDH servo (i.e., output noise / servo TF)? If so, it seems pretty high. With the right components, you should to be able to do ~100x better.

Quote:

Tara was able to tame the cavity servo loop. The third attachment is the result of several SR785 measurements of the error signal power spectrum. I converted this to a frequency noise power spectrum (fourth attachment) by extracting the voltage/frequency calibration factor from the error signal as follows.

 

 

  1132   Thu Mar 28 12:36:27 2013 ZachDailyProgressElectronics EquipmentFrequency readout noise of PDH loop

Is this really the input noise of the PDH servo (i.e., output noise / servo TF)? If so, it seems pretty high. With the right components, you should to be able to do ~100x better.

Quote:

Tara was able to tame the cavity servo loop. The third attachment is the result of several SR785 measurements of the error signal power spectrum. I converted this to a frequency noise power spectrum (fourth attachment) by extracting the voltage/frequency calibration factor from the error signal as follows.

 

  1131   Wed Mar 27 01:52:34 2013 taraNotesPMCPMC Longitudinal Testing and Modal Analysis

I use COMSOL to find the first longitudinal mode of a stainless steel PMC,  it is about 16 kHz. I'll find an analytical solution and compare them to make sure that the FEA result gives us a reasonable answer or not.

The FEA result in psl:1088 does not show the right body mode of the PMC. The frequency of 440 Hz is from some weird mode as seen from the figure in the entry. Evan checked the body mode of a simplified steel PMC, and I also check independently. Our results agree quite well that the first longitudinal mode is at ~16kHz.

pmc_eigen.png

However, this does not answer what we measured in PSL:1097, where the longitudinal motion is around 300Hz. I checked the body frequency of the base blocked and it is even higher than the PMC body modes' frequencies (this should be expected since the base is even bulkier).

Note: I just learned from Zach that the PMC in GYRO setup does not have 3-point support. It just sits on the base block. But this has not given me any clues about the possible modes yet.

 I'm writing some background and requirement for the PMC[coming soon]

 

  1130   Tue Mar 26 22:42:23 2013 EvanDailyProgressopticMode matching extravaganza

Tara and I repositioned the QWP and PBS immediately preceding the periscope so that we could move the 64.4-mm ROC lens closer to the cavity. For space reasons, this lens is now forked directly to the table rather than mounted on a translation stage. I tried for a while to adjust this and the 38.6-mm ROC lens to improve the mode matching, but I can't seem to do much better than 80% visibility. We may have to adjust the 103-mm ROC lens directly after the PMC in order to go further.

In better news, we were able to couple some power into the fiber that runs into the ATF. The beam is picked off with the PBS immediately following the EAOM and then sent through two mode-matching lenses and a HWP before hitting the fiber. We're sending 10 mW in and currently getting 0.85 mW out. More work is needed to get the polarization correct and to improve the coupling efficiency. This setup will probably have to be redone at some point, since the current pickoff beam is downstream of the cavity EOM and therefore has sidebands on it. Also, we will have to redo the coupling if we touch the 103-mm ROC lens to improve the cavity mode matching.

  1129   Tue Mar 26 10:00:19 2013 EvanDailyProgressElectronics EquipmentFrequency readout noise of PDH loop

Quote:

Today I tuned the mode matching into the south cavity by adjusting the two lens translation stages. The 64.4-mm ROC lens is at 15.63 mm, and the 38.6-mm ROC lens is at 4.01 mm. We currently have 80% visibility (see first attachment). One of the lenses is at the end of the range of its stage, so we will have to work to reposition it and continue mode matching.

Tara was able to tame the cavity servo loop. The third attachment is the result of several SR785 measurements of the error signal power spectrum. I converted this to a frequency noise power spectrum (fourth attachment) by extracting the voltage/frequency calibration factor from the error signal as follows.

I smoothed the PDH error signal by convolving with a gaussian and then numerically differentiated (second attachment). The locations of the carrier and the sidebands can then be read off, along with the magnitude of the derivative at zero cavity offset. Sidebands are at −14.5 ms and +14.7 ms, and the carrier is at 0.2 ms. Since the sidebands are 14.75 MHz from the carrier, this gives the time-voltage calibration as 1.01 GHz/s. The zero crossing of the carrier has derivative 16600 V/s, and hence the PDH slope is 61 kHz/V.

 I dialed the 38.6-mm ROC lens back to 1.11 m, twiddled the periscope mirrors, and got only marginal improvement in the visibility.

  1128   Mon Mar 25 22:56:42 2013 EvanDailyProgressElectronics EquipmentFrequency readout noise of PDH loop

Today I tuned the mode matching into the south cavity by adjusting the two lens translation stages. The 64.4-mm ROC lens is at 15.63 mm, and the 38.6-mm ROC lens is at 4.01 mm. We currently have 80% visibility (see first attachment). One of the lenses is at the end of the range of its stage, so we will have to work to reposition it and continue mode matching.

Tara was able to tame the cavity servo loop. The third attachment is the result of several SR785 measurements of the error signal power spectrum. I converted this to a frequency noise power spectrum (fourth attachment) by extracting the voltage/frequency calibration factor from the error signal as follows.

I smoothed the PDH error signal by convolving with a gaussian and then numerically differentiated (second attachment). The locations of the carrier and the sidebands can then be read off, along with the magnitude of the derivative at zero cavity offset. Sidebands are at −14.5 ms and +14.7 ms, and the carrier is at 0.2 ms. Since the sidebands are 14.75 MHz from the carrier, this gives the time-voltage calibration as 1.01 GHz/s. The zero crossing of the carrier has derivative 16600 V/s, and hence the PDH slope is 61 kHz/V.

Attachment 1: res_and_pdh.pdf
res_and_pdh.pdf
Attachment 2: errsig_smooth_derivative.pdf
errsig_smooth_derivative.pdf
Attachment 3: errsig_voltage_noise.pdf
errsig_voltage_noise.pdf
Attachment 4: errsig_freq_noise.pdf
errsig_freq_noise.pdf
  1127   Mon Mar 25 16:10:19 2013 EvanNotesPMCEigenfrequencies of PMC body: silica and stainless steel

Quote:

I asked Comsol for the eigenfrequencies of a simplified PMC body. The outer dimensions are as in the design document (6.89″ × 2.375″ × 2.00″), and the borehole has a uniform diameter of 1.188″ (instead of stepping down to a smaller diameter part-way through the body).

Comsol says the lowest mode for silica is at 8.3 kHz, and for stainless steel the lowest mode is at 7.2 kHz. For this simulation the body is assumed to be completely free; I didn't add 3-point contacts or anything like that.

 The lowest longitudinal mode for silica is 16.4 kHz, and for steel is 14.2 kHz.

  1126   Sat Mar 23 12:52:48 2013 EvanNotesPMCEigenfrequencies of PMC body: silica and stainless steel

I asked Comsol for the eigenfrequencies of a simplified PMC body. The outer dimensions are as in the design document (6.89″ × 2.375″ × 2.00″), and the borehole has a uniform diameter of 1.188″ (instead of stepping down to a smaller diameter part-way through the body).

Comsol says the lowest mode for silica is at 8.3 kHz, and for stainless steel the lowest mode is at 7.2 kHz. For this simulation the body is assumed to be completely free; I didn't add 3-point contacts or anything like that.

Attachment 1: pmc_block_sim.jpg
pmc_block_sim.jpg
  1125   Sat Mar 23 12:33:21 2013 EvanDailyProgressElectronics EquipmentError signal noise and stability

So far, we've been locking the south cavity by feeding back only on the laser frequency, not on the broadband EOM. This is because the demodulation path produces an error signal with the wrong sign for the EOM feedback, and the only way to flip the sign is to manually adjust cable lengths.

Last night I shortened the cable length between the LO 4-way splitter and the LO input to the frequency TTFSS in order to flip the sign of the error signal. The phasing is not quite optimal but it's pretty close (the cable length probably requires ~10 cm of adjustment at most). The power into the LO input was 5.9 dBm with the old cable, and is now 6.1 dBm with the new cable.

I then toggled the polarity of the inversion on the TTFSS from (+) to (−) and turned on the servo with laser frequency feedback only. I got the cavity to lock, but the error signal had huge 8 kHz sawtooth oscillations that I hadn't seen before (1st picture, below; trace 1 is cavity refl power, trace 2 is the error signal from OUT2 on the TTFSS). Then I added EOM feedback to the servo, and got huge 5 kHz sawtooth oscillations along with some spiky crap (2nd picture, below). Then I turned down the common gain knob to about 100 and the fast gain knob up to about 600, and got the loop to "stabilize" in the sense that the error signal now oscillates in the same way as before; i.e., with 200 kHz sinusoidal oscillations (3rd picture, below — note the similarity to the error signal oscillation in post 1123).

I took a quick spectrum of the error signal corresponding to the 3rd picture (attached below) with the SR785. Since the sinusoidal oscillations are happening at 200 kHz or so, this spectrum isn't capturing these oscillations.

Attachment 1: laser_only_bad.jpg
laser_only_bad.jpg
Attachment 2: laser_and_eom_bad.jpg
laser_and_eom_bad.jpg
Attachment 3: laser_and_eom_better.jpg
laser_and_eom_better.jpg
Attachment 4: laser_and_eom_better_spec.pdf
laser_and_eom_better_spec.pdf
  1124   Wed Mar 20 22:18:28 2013 taraNotesRefCavestimated beat frequency

At current temperature, the estimated beat frequency will be ~ 60-100MHz. This is not so bad, since we can use 1811 to measure the beat signal and use PLL to extract the beat noise.

 

We will need to use thermal expansion to tune the beat frequency. So, as a start, I try to figure out the beat frequency, and how much we have to heat up the cavity. The heaters on each cavity is off, only the heater around the vac chamber is on (but the servo is off). 

Right now we have one laser locked to one cavity, but the beam path to the first cavity has a beam splitter that we can borrow the beam and direct it to the 2nd cavity. I realigned the beam to have both beams into both cavities. By adjusting the temperature control on the NPRO (slow signal), I can bring the beam to resonant in each cavity.

1st cav is resonant @ (334/398) and (150/398). The numbers correspond to coarse and fine knobs of the slow feedback to the laser.

2nd cav is resonant @ (154/398). (I'll come up with a better name to call the cavities)

The FSR is 4.07 GHz (for 1.45" long cavity). This means 334-154 = 184 clicks on the coarse knob equals to 4.07 GHz, or 22MHz per coarse click. Both cavities resonant at ~3-5 clicks apart. So the beat frequency is ~ 60-100MHz. This is quite good, at least we are not close to half FSR apart. The power required to tune the cavity length should not be that high.

 

The next thing to do is try to see which cavity we need to heat up in order to bring both cavities resonant frequency closer together.

  1123   Wed Mar 20 20:44:32 2013 EvanDailyProgressopticAdded Faraday isolator

Today I installed the Faraday isolator after the PMC. Tara and I then spent some time trying to figure out why the PDH error signal suddenly had a huge DC offset (it was because I accidentally knocked the angle control on one of the HWP mounts while installing the FI beam dump). Before installing the FI, we had observed that the loop oscillates noticeably at about 100 kHz and had hoped it was caused by back-reflection into the laser (which the FI would fix). Installing the FI seems to have no effect on the oscillations. After installing the FI I adjusted the HWP immediately following and retuned the phasing of the PDH loop by adding some extra cable to the PD SMA input. I've attached a picture showing the sweeps of the cavity refl response and PDH error signal, and a picture showing the oscillations when the loop is engaged.

I tried minimizing the rejected light out of the FI to optimize the angle of the QWP directly in front of the cavity, but this light appears to be dominated by reflections other than those off of the cavity. The rejected light consists of two distinct spots which can be seen with an IR card. I think one of them is a reflection from the lens immediately following the FI, and the other is a reflection from the 14.75 MHz EOM.

I also tried improving the mode matching into the cavity by adjusting the two lenses on translation stages, but I didn't get much improvement. Right now we are at just under 50% coupling.
Attachment 1: cavity_scan.jpg
cavity_scan.jpg
Attachment 2: loop_oscillation.jpg
loop_oscillation.jpg
  1122   Sun Mar 17 16:09:10 2013 EvanNotesopticAlAs/GaAs

We should be able to mode match into a cavity with 1.0 m ROC mirrors using only the optics we already have on the table. 

Current mirrors: 0.5 m ROC (has -1114 mm FL)

  • 370 um PMC waist at z = 0 m = 0 in
  • 229.1 mm FL lens at z = 0.203 m = 8.0 in
  • 209 um intermediate waist
  • 85.8 mm FL lens at z = 0.923 m = 36.3 in
  • 41 um intermediate waist
  • 143.2 mm FL lens at z = 1.201 m = 47.3 in
  • 182 um cavity waist at z = 1.944 m = 76.5 in
  • Mode overlap 1.000

Proposed mirrors: 1 m ROC (has -2227 mm FL)

  • 370 um PMC waist at z = 0 m = 0 in
  • 229.1 mm FL lens at z = 0.203 m = 8.0 in
  • 209 um intermediate waist
  • 85.8 mm FL lens at z = 0.889 m = 35.0 in
  • 45 um intermediate waist
  • 143.2 mm FL lens at z = 1.166 m = 45.9 in
  • 210 um cavity waist at z = 1.952 m = 76.8 in
  • Mode overlap 0.999

The various waists for the proposed mode matching are equal to or larger than the waists for the current mode matching, so I don't think we should be any more worried about sensitivity than we already are.

 

Quote:

I'm thinking about the spec for AlAs/GaAs coatings. Here is the list of what I have:

  • coating on concave side of the mirror for 0.5m x6 (I'm not sure if they can do the transfer on 0.5m mirror now) for 1.0m x6 for flat mirror x3 -
  • for circularly polarized light, normal incidence
  • Transmission @1064 = 100ppm +/- 10ppm. 10% error is still within the acceptable value for 10ppm loss (T ~ 67-73%), see T1200057v11 -
  • Absorption + scatter loss < 10ppm, this is what Garrett told us. -
  • coatings diameter = 8mm (The number is from Garrett), the loss around the edge for our beam with diameter=364 um is less than 10^-10 ppm. -
  • Max scratch surface and point defects are not determined yet. I can look up the specs from our current SiO2/Ta2O5 mirror since they are ok for us. -
  • I think we are aiming for the thermo-optic optimized coatings. The layer structure can be found in T1200003-v1.

==Coating diamter for 0.5m ROC mirror==

About the coatings diameter, Garrett said it depends on the aperture size/ coating diameter. So I made a plot to estimate the loss due to the finite size coating vs Coating diameter for our spot radius of 182 um. The loss is simply calculated by the ratio of the power not falling on the coating = Ploss/Pin = (exp(-2*r0.^2./w0.^2))*1e6*26000/pi   

where r0 = coating radius, w0 = spot radius, a factor of 1e6 for showing the result in ppm, 26000/pi is the total loss due to the light bouncing in the cavity.

edgeloss.png

fig1: Loss vs coating diameter (in meter)

 

It seems we can go to 2mm coating diameter, and the loss is still much less than 1ppm (the expected loss from absorption and scatter is ~ 10ppm). However, we have to consider about how well they can center the film, how well we can assemble the cavity. So larger coating diameter is always better. If we assume that 1mm error is limiting us, coating diameter of 4-5 mm should be ok for us.

 ==for mirror with 1m ROC==

If the ROC is 1.0m, the coating diameter can be 8mm. For the cavity with 1.45" long, the spot radius on the mirror will be 215um (182um with 0.5m mirror). This changes the noise budget of the setup a little bit. The total noise level is lower by a factor of ~ 1.2. (see below figure) at 100 Hz.

noise_compare.png

fig2: Noise budget comparison between setup with 0.5 m and 1.0m RoC mirrors, plotted on top of each other. Noises that change with spotsize are coating brownian, substrate brownian, thermoelastic in substrate, and thermo-optic.

 

==What do we choose? 0.5m or 1.0m==

For both 0.5 and 1m, the cavity will be stable (see T1200057-v11, fig11). So either choice is fine

if we use 1.0 m,

  • we loss the signal level a bit,
  • but we are more certain that the coating will work. 
  • The procurement should be faster (as promised by Garrett)
  • have large area coating up to 8mm diamter
  • need to check if we can mode match or not (I'm positive that we can, but I'll check or let Evan check)

So at this point, I'm thinking about going with 1.0 m mirror.

 

 

 

Attachment 1: 0.5m_ROC_modematch.pdf
0.5m_ROC_modematch.pdf
Attachment 2: 1m_ROC_modematch.pdf
1m_ROC_modematch.pdf
  1121   Fri Mar 15 11:23:45 2013 taraNotesopticAlAs/GaAs

I'm thinking about the spec for AlAs/GaAs coatings. Here is the list of what I have:

  • coating on concave side of the mirror for 0.5m x6 (I'm not sure if they can do the transfer on 0.5m mirror now) for 1.0m x6 for flat mirror x3 -
  • for circularly polarized light, normal incidence
  • Transmission @1064 = 100ppm +/- 10ppm. 10% error is still within the acceptable value for 10ppm loss (T ~ 67-73%), see T1200057v11 -
  • Absorption + scatter loss < 10ppm, this is what Garrett told us. -
  • coatings diameter = 8mm (The number is from Garrett), the loss around the edge for our beam with diameter=364 um is less than 10^-10 ppm. -
  • Max scratch surface and point defects are not determined yet. I can look up the specs from our current SiO2/Ta2O5 mirror since they are ok for us. -
  • I think we are aiming for the thermo-optic optimized coatings. The layer structure can be found in T1200003-v1.

==Coating diamter for 0.5m ROC mirror==

About the coatings diameter, Garrett said it depends on the aperture size/ coating diameter. So I made a plot to estimate the loss due to the finite size coating vs Coating diameter for our spot radius of 182 um. The loss is simply calculated by the ratio of the power not falling on the coating = Ploss/Pin = (exp(-2*r0.^2./w0.^2))*1e6*26000/pi   

where r0 = coating radius, w0 = spot radius, a factor of 1e6 for showing the result in ppm, 26000/pi is the total loss due to the light bouncing in the cavity.

edgeloss.png

fig1: Loss vs coating diameter (in meter)

 

It seems we can go to 2mm coating diameter, and the loss is still much less than 1ppm (the expected loss from absorption and scatter is ~ 10ppm). However, we have to consider about how well they can center the film, how well we can assemble the cavity. So larger coating diameter is always better. If we assume that 1mm error is limiting us, coating diameter of 4-5 mm should be ok for us.

 ==for mirror with 1m ROC==

If the ROC is 1.0m, the coating diameter can be 8mm. For the cavity with 1.45" long, the spot radius on the mirror will be 215um (182um with 0.5m mirror). This changes the noise budget of the setup a little bit. The total noise level is lower by a factor of ~ 1.2. (see below figure) at 100 Hz.

noise_compare.png

fig2: Noise budget comparison between setup with 0.5 m and 1.0m RoC mirrors, plotted on top of each other. Noises that change with spotsize are coating brownian, substrate brownian, thermoelastic in substrate, and thermo-optic.

 

==What do we choose? 0.5m or 1.0m==

For both 0.5 and 1m, the cavity will be stable (see T1200057-v11, fig11). So either choice is fine

if we use 1.0 m,

  • we loss the signal level a bit,
  • but we are more certain that the coating will work. 
  • The procurement should be faster (as promised by Garrett)
  • have large area coating up to 8mm diamter
  • need to check if we can mode match or not (I'm positive that we can, but I'll check or let Evan check)

So at this point, I'm thinking about going with 1.0 m mirror.

 

 

  1120   Thu Mar 14 22:39:49 2013 EvanDailyProgressRefCavSouth cavity locked on TEM00

 [Tara, Evan]

Tara and I spent today and Tuesday laying down the new optical path for the south cavity according to the layout in entry 1100. After quite a bit of periscope steering (mostly on Tara's part), we got flashes of TEM00 and subsequently were able to lock the laser's fast frequency control servo to the cavity resonance. Getting near the resonance requires a bit of dialing around with the set point on laser temperature servo; the values that worked today were 512 on the coarse adjust and 894 on the fine adjust.

From here, we should

  1. make a new mount for the post-PMC Faraday isolator, since the existing mount is about 32 mils too high (and hence the FI has not yet been placed in the beam path);
  2. fine-tune the steering into the cavity;
  3. optimize the mode-matching into the cavity;
  4. measure the cavity finesse;
  5. pick a different lens for matching into the RFPD, since the current RFPD path length is too long; and
  6. fine-tune the demodulation phase for the loop.
  1119   Thu Mar 14 18:20:04 2013 taraNotesopticfused silica pmc

Peter told me that  the fused silica pmc currently used in the lab is bonded by Vac-seal epoxy. So we don't need to polish any surfaces for optical contact.

IMG_0819.JPG

Traces of vac-seal can be seen between the mirror and the tip, the tip and the spacer bonded areas. Vac-Seal epoxy is chosen for its low out gasing, so that the mirrors won't be contaminated.

IMG_0820.JPG

  1118   Thu Mar 14 17:27:42 2013 KojiDailyProgressVacuumCavities inserted; chamber pumping down

Slight baking with the heaters? This also gives you the test of the heater and temp sensors.

  1117   Wed Mar 13 23:52:12 2013 taraDailyProgressVacuumCavities inserted; chamber pumping down

It's still improving, now the current is 0.3 mA (0.38mA yesterday). I'll wait until it stops improving and try to tighten the screws a bit. \

About the plastic pieces, they are peek. I think it is vacuum compatible, cf E960050-v11.

May be it is the temp sensor that I have to re-solder on the copper pieces. I did not bake them after soldering.

Quote:

The gap at the frange is OK as long as the gasket is evenly squeezed.

1. It the pressure still improving? => Your cavity is still out gassing.

2. Fasten the screws little more (not too much) for these two franges.
    => If there is any improvement there is some leakage.
    => If not, this is just a outgas from your cavity.

Do you think your black plastic pieces are vacuum compatible? Are they made of Delrin?

 

  1116   Wed Mar 13 10:50:31 2013 KojiDailyProgressVacuumCavities inserted; chamber pumping down

The gap at the frange is OK as long as the gasket is evenly squeezed.

1. It the pressure still improving? => Your cavity is still out gassing.

2. Fasten the screws little more (not too much) for these two franges.
    => If there is any improvement there is some leakage.
    => If not, this is just a outgas from your cavity.

Do you think your black plastic pieces are vacuum compatible? Are they made of Delrin?

  1115   Tue Mar 12 20:51:00 2013 ranaNotesBEATMode matching plan for beat breadboard

 

 Its good that the beams can be matched, but there are a few more considerations to take into account:

* How does this optimize the geometry to make the readout more insensitive to vibrations? Will it all fit inside a well insulated plastic box?

* How does the final beam size depend on the physical parameters (i.e. path length, lens Radius error, etc.)? i.e. we need a sensitivity analysis.

* The divergence angle of the beam at the detector determines, in part, how sensitive this setup is to backscatter from the RFPD. How to minimize this? How are the specular reflections from the PD face dumped?

  1114   Tue Mar 12 18:20:21 2013 taraDailyProgressVacuumCavities inserted; chamber pumping down

After pumping the chamber down for two days, I disabled the turbo pump and turned on the ion pump. The initial current was 7mA. After a day now it is 0.38 mA. It was better before, see PSL842, (started the ionpump at 1mA, and operated at less than 0.1 uA). If this is true, this means the pressure will be about a factor of 0.38mA/0.1uA ~ 3800 higher than before. (the calculation is based on this equation, where the current is directly proportional to the pressure.

 

The flanges do not flatly touch each other. There's a tiny bit of something in the middle that causes a gap as shown below. This might be the reason why the ion pump current is high, or it might be some out-gasing problems of the in-vac materials. I think we can try to spray isopropanol around the flange to see if the current comes up, but I think it is a bad idea for the ion pump. I'll ask Steve or Koji for their opinions.

 IMG_0815.JPG

The pictures during the installation can be found onpicasapage

Quote:

 [Tara, Koji, Evan]

Tara and Koji spent the better part of yesterday afternoon inserting the new cavity assembly into the vacuum chamber. In the process of putting the window back on the chamber, the old copper gasket may or may not have hit the inner surface of the window, so Koji performed a drag wipe. Tara and Koji then inspected the window under a high-power bulb, and I think the consensus is that there's no visible damage.

The chamber is currently pumping down. Unfortunately, there appears to be a small speck of something trapped between the window and chamber flanges, so there's a small gap on one side of the joint. We'll see if we can achieve high vacuum.

 

  1113   Sun Mar 10 14:23:56 2013 EvanNotesDrawingsMode matching plan for beat breadboard

I ran CVI's list of 1064-nm silica lenses through a la mode looking for a good way to get the cavity beams onto the beat PD. We're looking for a spot with radius somewhere between 80 and 150 um at the PD.

I tried a few optimizations with only one lens in each beam's path, but the results weren't very good; the beam is either too large, too small, or requires excessive path length after the BS. What I've attached is the result of an optimization for two lenses in each beam's path. It gives a 102 um waist a few inches after the last folding mirror, which is nice. The spot radius at the BS is 470 um, which is not too large.

Attachment 1: beat_modematch_layout.pdf
beat_modematch_layout.pdf
Attachment 2: beat_modematch_alm.pdf
beat_modematch_alm.pdf
  1112   Sat Mar 9 11:38:10 2013 EvanDailyProgressVacuumCavities inserted; chamber pumping down

 [Tara, Koji, Evan]

Tara and Koji spent the better part of yesterday afternoon inserting the new cavity assembly into the vacuum chamber. In the process of putting the window back on the chamber, the old copper gasket may or may not have hit the inner surface of the window, so Koji performed a drag wipe. Tara and Koji then inspected the window under a high-power bulb, and I think the consensus is that there's no visible damage.

The chamber is currently pumping down. Unfortunately, there appears to be a small speck of something trapped between the window and chamber flanges, so there's a small gap on one side of the joint. We'll see if we can achieve high vacuum.

  1111   Wed Mar 6 21:56:12 2013 taraDailyProgressVacuumnote on feedthrough

Feedthrough channel (as seen from the connector outside of the chamber)

1-6: heater on cavity#98 : 85.4 Ohm

3-7: Temp sensor on cavity #98

 

4-8: heater on cavity#99: 156 Ohtm

5-9: Temp sensor on cavity#99

 

  1110   Wed Mar 6 15:44:26 2013 ZachDailyProgressTempCtrlHeater for refcav

Is there a reason you don't want to use a Daedalus board (see bunch of ATF entries here). There will need to be some modifications, but besides these minor things it sounds like exactly what you want.

The only major thing is that I stupidly picked the SO8 package version of the BUF634, which can only dissipate less than a watt. Using one of these nifty SO8-to-DIP converters, you can replace it with a TO-220 package and have several watts.

A particularly nice thing is that each board has two channels, if you're into that sort of thing.

Anyway, let me know. I can help you get started when I'm at Caltech later this month if you decide you want it. I need to finish setting up the one for the gyro, anyway.

Quote:

 To make a low noise current driver, use the BUF634 wrapped around the loop of a low noise opamp (never, ever use the AD797) as illustrated in the BUF634 datasheet.

Don't use a commercial power supply, make a protoboard setup for a NIM box.

 

  1109   Wed Mar 6 12:48:50 2013 taraDailyProgressVacuumthermal shield + peek supports are baked

I setup the small vacuum chamber to bake the shields and peek cavity supporting pieces. All pieces are baked and I'm assembling all the parts.

 photo(13).JPG

Details about how to use the pump and the chamber can be found in CTNwiki.

  1108   Sun Mar 3 22:08:26 2013 taraDailyProgressVacuumopening vac chamber

 I found all necessary parts for the heater (crimp connectors, heating wire). I'll bake all parts once all the wiring is ready

 

 IMG_2255.JPGIMG_2258.JPG

fig1: left Dsub connector for the cable, Right, heater(yellowish wire around the tube) is connected to wiring cable with crimp connectors.

download.jpg

fig2: crimp connector( vacuum compatible material). I need to borrow the suitable crimp tool from Down, see PSL:775

heater.png

fig3: above, previous wiring (one heater,3 sensors), below current wiring (two heaters, two sensors)

  1107   Thu Feb 28 09:12:22 2013 not taraDailyProgressTempCtrlHeater for refcav

 To make a low noise current driver, use the BUF634 wrapped around the loop of a low noise opamp (never, ever use the AD797) as illustrated in the BUF634 datasheet.

Don't use a commercial power supply, make a protoboard setup for a NIM box.

  1106   Wed Feb 27 10:16:09 2013 koji, taraDailyProgressVacuumopening vac chamber

We opened the vacuum chamber and brought the stack with the 8" cavities out to the clean bench. New 1.45" cavities are under preparation to go in the chamber.

The 8" cavities and the double seismic isolation stack were removed from the chamber. The connector had to be removed from the inside of the mini flange for the feedthrough [add pic]. We replaced the top seismic stack along with 8" cavities and their mounts, with the new stack/new mount for the shorter cavities. We reuse the bottom stack.

Next is to modify the wiring for heaters and temp sensors. Currently, the connector is wired for 3 sensors and 1 heaters (for 9-pin connector). Soldering at the 9-pin connector seems to be a tough job. Koji suggested that I remove one of the temp sensor at its legs, not at the connector end. Then connect the unoccupied cables to the heating wire.

  1105   Wed Feb 27 09:57:33 2013 taraDailyProgressTempCtrlHeater for refcav

I'm contemplating about how to drive the heater of the copper shields. So here is a list of what happening.

Objective: to drive the heater with ~ 1W power (see the previous entry). The estimated power is from my calculation. I have not taken heat conduction/uncertainties of surface emissivity into account, we might need more power to be on the safe side.

Problem: If we use a low noise driver (using AD797, cf PSL:765), the maximum delivered power will be only~0.5 W.

So, do we need the low noise driver? or do we need to have enough power? If we just need enough power, we can drive the heater with a commercial power supply.

At this point, I think it is more important to be able to tune the cavity length with larger range so that we can see the beat. Thus, sufficient heating power should be our top priority . Better heater/ control can be developed next, once we see the beat.

So the plan is:

  1. Wrap both shields up with heating wire. The length of the wire (total resistance) depends on the limiting current/voltage available from a power supply to ensure ~2W power.
  2. Put the cavity back and bring the laser to the cavities.
  1104   Mon Feb 25 01:02:58 2013 taraDailyProgressVacuumHeater for refcav

I did the calculation to estimate the required power for heating up the cavity by 20 K above room temperature. More details are coming.

 ==Motivation==

Since we plan to tune the beat frequency by tuning the cavity length with thermal expansion, we need to know how much power is needed to heat one cavity up by 20K above room temp (see LIGO-1200057 for more detail).

 ==Model==

To simplify the calculation, I use 2-D model for the cavity and the shield. Assuming the system is in equilibrium. In the calculation, I considered the effect from reflected radiation from both the shield, and the cavity itself.

photo(10).JPG

 ==result==

  •   At room temperature of 29 C, depending on the emissivity of the shield, the tube has to be from 20 C - 100 C above room temperature inorder to keep the cavity at 20 C above room temp. Since the emissivity of the copper can vary from, .03 (highly polished)-0.77 (oxide),  see ref. Our cavity is polished a bit, so the reasonable value should be ~ 0.1-0.2.  Thus, the shield temperature should be able to reach at least 50 [C]. 

heater.png

 

 

  •  Next is to estimate how long the wire should be for bringing the shield up to 50 degree C. If all the heating power from the wire goes to radiation, the radiated power from 50[C] shield, at room temp is sigma*epsilon*area*(T^4 - T0^4) which is ~ 0.6 - 1 Watt Note: I checked the calculation and Frank's note in PSL:765, the calculated number is reasonable.
  1103   Sun Feb 24 01:53:43 2013 taraNotesVacuumforgot to turn off the ion pump while venting

I vented the vacuum chamber and forgot to turn of the ion pump. It was on for ~ 5-10 minutes. I'm not sure if it is broken or not. I talked to Alastair and he said the metal plate might be contaminated and needed to be repair. I'll keep this in mind and see what happen when I pump down the chamber.

  1102   Sun Feb 24 01:16:26 2013 taraDailyProgressRefCavrefcav and mount assembly

Thermal shields and caps are ready. I cleaned them in ultra sonic bath with Isopropanol. They fit nicely to the mount.

The caps are teflon.

photo(9).JPG    photo(8).JPG

 

==Wire Heater==

I'm think about how to wire the heater around the cavity. I'm reading Frank's entries in PSL:765,768 ,776,786, . Seems like I might need to drill a few holes on the shields for the wires.  There are still more similar wire left. I'll calculate how long the wire has to be for heating the cavity up by 20 K.

 

  1101   Mon Feb 18 03:38:20 2013 taraNotesScheduleTask Sheet

Now that the layout is ok and the table is cleared, the next thing to do is open the chamber. The plan for this week will be:

==Cavities & Vac chamber==

  1. clear the clean bench
  2. Prepare all parts  for the new cavities that will go to the chamber. The assembly for the new cavities is almost done, I'm wating for the caps of the thermal shield.
  3. Open the chamber, remove the 8" cavities.
  4. Change the heaters/ thermometers. Currently, the feedthrough for the chamber is 9-pin. It is for 3 thermometer and 1 heater. The new setup will need 2 thermometers, 2 heaters. I don't know what kind of wire and where it is in the lab yet. This need to be checked.
  5. Once the sensors/heaters are installed, we can close the chamber and start the mode matching.

==1st laser==

  1.  The setup for the first laser should be easy. The setup from the laser to the PMC is not changed much except the BB eom for locking. I should be able to lock the first cavity soon

==2nd laser==

  1. Need to think about the PMC
  2. The EOM
  1100   Mon Feb 18 03:16:44 2013 taraNotesDrawingslayout for the new setup

The new optic layout is done. There are some changes from the previous version.

Some notes and concerns about the new layout:

  • I made sure that the beam size in all EOMs, EAOMs are in the right size (diameter ~ 250 - 500 um) to avoid any clipping/ diffraction.
  • Due to the small waist size of the new cavities and the long path between the cavity and the vacuum window. The beamsize on the input optic is quite large (almost 3mm). I'll have to check if the beam spiltter is large enough.
  • I assume that the beam size from the new PMC (the setup on the left) is 370um. It might change because the design for the PMC is not settled yet, but the spacer is enough for mode matching.
  • I add a beam splitter in the setup just before the vacuum window. This can be used for RFAM pickup or fiber distribution. It can be done in the future.
  • a set of flip mirrors are added, for scanning two cavities with one laser. The mode match should be good enough to get TEM00 in the 2nd cavity.

ctn_layout_2laser_v2.png

 

  1099   Wed Feb 13 19:54:32 2013 taraNotesLaserexternal cavity laser diode

Some notes about external cavity diode laser. I investigated in this about a year ago.  I think it might be a good time to work on it, since I'm modifying the  ctn layout and we can use a frequency stabilized laser (although locked to a short cavity) via fiber optic to test with a home made ecdl. 

Quote:

I just asked Aki a few more questions about the ecd laser. If we do not require the performance to be rival to that of the NPRO, making one is possible in a few weeks time scale.

Q1) The setup of the Littrow style laser you showed me had one mirror
behind the grating. Is the setup similar to this
<
http://rsi.aip.org/resource/1/rsinak/v72/i12/p4477_s1>? Where the
mirror is used so that the alignment does not change when the laser is
tuned.

A1) Actually, Yes, for the laser you took the photo. But, most of the lasers we use in the lab don't have a mirror, only diode and grating in the box. In our case, we only scan the frequency by ~1GHz and the pointing vector drift is negligible.

 

Q2) Did you remove the glass window of the diode laser when you assemble
 the laser? If so, how do you keep the diode clean, or it does not matter
for your requirement.

A2) We didn't remove the glass window. What we did is very simple. We mounted the bare diode on the thorlabs mount: http://www.thorlabs.com/thorProduct.cfm?partNumber=LT230220P-B

 

 Q3) You mentioned that the line width of the laser when locked to CS cell
is ~300kHz. Is it because of gain limited of the servo or the CS cell's
intrinsic noise?

A3) We haven't measure the linewidth with locking and without locking independently. It's possible that our laser linewidth (without frequency lock) might be ~300kHz. So I don't know what limits our linewidth.

One thing you may want to consider is that a diode laser is infamous for the broad background incoherent light, compared to the solid state lasers. We typically observe ~30nm-wide incoherent light around the carrier with 30-40dB suppression compared to the carrier. If your experiment is sensitive to the spectral purity, this might be an issue.

Aki

 

 So the question is do we want to try to build one similar to what they have? We know that with the time scale and experience we have it will not be as good as the performance of the ecd laser reported in Numata etal paper, but it might be a fun project for the SURF student.

Quote:

I'm looking into how to make external cavity diode laser (ecd).

Here is the list of what we need.

 

  • Laser diode 1064nm, diode (+AR coating)
  • Grating
  • Current driver for locking the laser (home made)

I ask Akihisa who works in Kimble's group about their home made 850 nm ecl. The performance is not as good as NPRO yet (300kHz line width when locked to CS cell), but it is certainly interesting.

  • They use Littrow configuration ( 1 grating to form a cavity). The mirror behind the grating is for adjusting the output beam.
  • The PD is not AR coated, a slight power (10 uW) that transmits inside the diode is enough for the feedback.
  • The linewidth is ~ 300kHz when locked to CS cell, the free running noise is not measured.
  • They use both PZT and current to actuate on the frequency stabilization, the bandwidth is 50kHz.
  • Temperature feedback is employed to keep the stability of the laser
  • Time for putting everything in the box (once all components are ready) is ~ 1wk

 (IMG_0887.jpg IMG_0888.jpg

 

 

  1098   Fri Feb 1 03:14:45 2013 taraNotesDrawingslayout for the new setup

Quote:

I'm making a new layout for the 2 laser setup. It is in progress. I just want to make a note about some features I want to have in the setup.

  • Move EAOM (for intensity stabilization,ISS) some where down stream, close to the chamber. Since the monitor is behind the cavity.
  • Pick up for optical fiber: This should be somewhere behind 1)EAOM, so that power can be stabilized, 2) BB EOM, so that the laser frequency is stabilized at higher frequency as well. 
  • Some beam splitter/ flip mirrors to bring laser to both cavities. so that I can scan two cavities with 1 laser at the same time. This should be handy when we need to find the beat, since the beat frequency is not well determined (by AOM) anymore.
  • Some space for an independent frequency stabilized laser. This is for the future, when we want to lock a laser to the 8" cavity and distribute it to other labs.  From the current setup, I think we might have to borrow the laser from the 2-laser setup, after the PMC, then lock it to the 8" cavity. Installing 3 sets of laser, PMC, refcav on the table is not easy. With this idea, we can lock the laser to a low noise cavity(8"), and just change the cable, when we want to switch to do the coating measurement.

 The tentative plan is shown below. I also move the BB eom for locking frequency(phase) next to the laser.

  • The pick up for fiber distribution can be at the PBS just after the PMC where we dump the power. However the laser is not intensity stabilized yet, but I'll worry about it later.
  • I add pick up for RFAM monitor. It is the unused beam after the EAOM. The power will fluctuate when ISS is active, but should be fine for RFAM monitoring purpose.

I'm doing mode matching calculation to see if the lenses can be placed in the available area. Otherwise I need to fold the beam a few more times.

I also started removing the acoustic box, some optics on the table so that I can try to position the real optics to see if they can fit on the table.

 

ctn_layout_2laser.pdf

  1097   Thu Jan 31 14:00:00 2013 Kristen HoltzNotesPMCPMC Longitudinal Testing and Modal Analysis

Longitudinal Mode frequency @ 420 Hz

 

Monday, January 28, 2013

Tests of resonant frequency specifically for longitudinal mode done on PMC because the piezo will only affect this length.  Originally the PMC spacer was constrained in motion by clamps on the PMC base, as shown below.

PMCphoto_circclamps.JPG

 Data was taken with an oscilloscope and laser vibrometer - the B&K system was not functioning correctly - and the PMC spacer was excited by hitting it with a marker.  The PMC was hit near the endcap, along the long axis of the PMC, as shown in the below image.

PMC_Excitation_Diagram_1-28-13.JPG

 

The results in the time domain are shown below and indicate a resonant frequency of 450 Hz.  A next step is to fit this data to a decaying sine wave.

PMC_ringdown_Jan28.JPG

When the above clamps were loosened to allow for more motion, the frequency dropped to closer to 350 Hz.

PMC_unconstrainedringdown_Jan28.JPG

 

ELOG V3.1.3-